find the point of intersction forthe line below using elimation 2x+3y=6, -2x+y=10
Answer by stanbon(75887)

find the point of intersction forthe line below using elimation
2x+3y=6,
-2x+y=10
Add the two equations to get:
4y=16
y=4
Substitute into 2nd equation to get:
-2x+4=10
-2x=6
x=-3
----------
Solution:
x=-3 ; y=4
Cheers,
Stan H.